# The Riemann Hypothesis is false

### v47

@@
$J(T):=\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right).$
Let $\phi(r)=\frac{T}{2\pi}\log(1+\frac{r}{m})$ thus $\phi^{\prime}(r)=\frac{T}{2\pi(r+m)}.$ Let $\eta\in(0,1/2)$ be a constant, $\beta=\phi^{\prime}(T^{1/4})$ and $\alpha=\phi^{\prime}(m\varepsilon)$. By Lemma 4.10 of Titchmarsh [7, p.76], we have
@@
Since $\beta-\alpha\ll\varepsilon\log^{2}T=o(1)$, note that the range $\alpha-\eta,\beta+\eta]$ contains at most one integer $v$. Hence (18) implies that
$J(T)\ll\Bigg|\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x\Bigg|+T^{-1/4}\log T.$ (19)
Define
$J_{0}(T)=\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi i(\phi(x)-vx)}\mathrm{d}x=\int_{T^{1/4}}^{m\varepsilon}\frac{1}{x}e^{2\pi iT\Big[\log(1+\frac{x}{m})-\frac{vx}{T}\Big]}\mathrm{d}x.\\$ (20)
Let $G(x)=\frac{1}{x},F(x)=\log(1+\frac{x}{m})-\frac{vx}{T},\lambda=2\pi T$ and $F^{\prime}(x_{0})=0$ so that $\frac{T}{m+x_{0}}=v$ thus $F^{\prime\prime}(x_{0})=-\frac{v^{2}}{T}$ and $|T-mv|\sim\frac{Tx_{0}}{m}$ where $x_{0}\in[T^{1/4},m\varepsilon]$. Then $J_{0}(T)$ is an oscillatory integral of the form $\int_{a}^{b}G(x)\exp(i\lambda F(x))\mathrm{d}x$. Thus by the stationary phase approximation 222see e.g. https://grokipedia.com/page/Stationary phase approximation (the one-dimensional section)., we have
$J_{0}(T)\sim G(x_{0})e^{i\lambda F(x_{0})-i\pi/4}\sqrt{\frac{2\pi}{\lambda|F^{\prime\prime}(x_{0})|}}\ll G(x_{0})\sqrt{\frac{2\pi}{\lambda|F^{\prime\prime}(x_{0})|}}\ll\frac{1}{|T-mv|}\ll T^{-1/4}$ (21)
uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ as $\lambda=2\pi T\rightarrow\infty$. Combing (19), (20) and (21) yields
$\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right)\ll T^{-1/4}\log T$ (22)
uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. Applying (22) into (17) gives
$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log^{2}T).$ (23)
Putting $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O({Tr^{2}/m^{2}})}=1+O({Tr^{2}/m^{2}})$ into the right-hand side of (23) gives
$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{2}T).$ (24)
Summing (24) over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\ \leq T(1+\varepsilon)$ yields
@@
as $X\rightarrow\infty$. 333see e.g. https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
@@
For $\|x_{m}\|>Z^{-1}$ we use $\frac{1}{1+Z\|x_{m}\|}\leq\frac{1}{Z\|x_{m}\|}.$ Partition $(0,1/2]$ dyadically: $2^{-j-1}<\|x_{m}\|\leq 2^{-j},0\leq j\leq\log_{2}Z.$ For such $m$, $\frac{1}{1+Z\|x_{m}\|}\ll\frac{2^{j}}{Z}$. Since $x_{m}$ is monotone in $m$, the number of $m$ in each dyadic interval is $\ll T\cdot 2^{-j}.$ Hence the contribution of each $j$ is $\ll T\cdot 2^{-j}\cdot\frac{2^{j}}{Z}=\frac{T}{Z}.$ Summing over $0\leq j\leq\log_{2}Z$ yields

### v48

@@
$J(T):=\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right)$ (18)
where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. Let $\phi(r)=\frac{T}{2\pi}\log(1+\frac{r}{m})$ thus $\phi^{\prime}(r)=\frac{T}{2\pi(r+m)}.$ Let $\eta\in(0,1/2)$ be a constant, $\beta=\phi^{\prime}(T^{1/4})$ and $\alpha=\phi^{\prime}(m\varepsilon)$. By Lemma 4.10 of Titchmarsh [7, p.76], we have
@@
Since $\beta-\alpha\ll\varepsilon\log^{2}T=o(1)$, note that the range $\alpha-\eta,\beta+\eta]$ contains at most one integer $v$. Summing (19) over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ gives
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}J(T)\ll\int_{T^{1/4}}^{T(1+\varepsilon)\varepsilon}\frac{1}{x}\Bigg|\sum_{m=\max\left(\frac{T}{(1+\varepsilon)\log^{2}T},\frac{x}{\varepsilon}\right)}^{T(1+\varepsilon)}e^{2\pi i\big(\phi(x)-vx\big)}\Bigg|\,dx\;+\;T^{3/4}\log T,$ (20)
where $v=v(m)$ is the integer nearest to $\frac{T}{2\pi m}$. The dependence of $v$ on $m$ must be respected, so we write $k=v(m)$. The condition $|k-\frac{T}{2\pi m}|\leq\frac{1}{2}$ implies $m\in I_{k}:=\Bigl[\frac{T}{2\pi(k+\frac{1}{2})},\frac{T}{2\pi(k-\frac{1}{2})}\Bigr].$ For $m\geq T/((1+\varepsilon)\log^{2}T)$ we have $0\leq k\ll\log^{2}T$; hence $k$ runs over $O(\log^{2}T)$ nonnegative integers. The length of $I_{k}$ is $|I_{k}|\asymp T(k+1)^{-2}$. On $I_{k}$ the phase function is $g_{k}(m)=\frac{T}{2\pi}\log\!\Bigl(1+\frac{x}{m}\Bigr)-kx$ where $k$ is fixed. Its derivative satisfies $g_{k}^{\prime}(m)=-\frac{Tx}{2\pi m(m+x)}\asymp-\frac{Tx}{m^{2}}$ and $|g_{k}^{\prime\prime}(m)|\asymp\frac{Tx}{m^{3}}.$ Applying the first‑derivative bound of van der Corput (Lemma 4.19 of Titchmarsh [7]) on each interval $I_{k}$ yields
$\Bigl|\sum_{m\in I_{k}}e^{2\pi ig_{k}(m)}\Bigr|\ll\frac{1}{\lambda_{1}}+|I_{k}|\lambda_{1},\qquad\lambda_{1}:=\frac{Tx}{m^{2}}\asymp\frac{x(k+1)^{2}}{T}.$
Since $|I_{k}|\asymp\frac{T}{(1+k)^{2}}$, we have
$\Bigl|\sum_{m\in I_{k}}e^{2\pi ig_{k}(m)}\Bigr|\ll\frac{T}{x}+x$
Summing over all admissible $k$ where $0\leq k\ll\log^{2}T$, gives
$\Bigl|\sum_{m}e^{2\pi i(\phi(x)-\nu x)}\Bigr|\ll\sum_{k}\Bigl(\frac{T}{x}+x\Bigr)\ll\frac{T\log^{2}T}{x}+x\log^{2}T.$ (21)
Inserting this bound into the integral in (20) gives
$\int_{T^{1/4}}^{T(1+\varepsilon)\varepsilon}\frac{1}{x}\Bigl|\sum_{m}e^{2\pi i(\phi(x)-\nu x)}\Bigr|\,dx\ll\int_{T^{1/4}}^{T^{3/4}}\Bigl(\frac{T\log^{2}T}{x^{2}}+\log^{2}T\Bigr)dx\ll T^{3/4}\log^{2}T.$
Plugging this into (20) gives
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}J(T)\ll T^{3/4}\log^{2}T.$ (22)
Combining (20), (21) and (22) yields
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{T^{1/4}\leq r\leq m\varepsilon}\frac{1}{r}\exp\!\left(iT\log\left(1+\frac{r}{m}\right)\right)\ll T^{3/4}\log^{2}T.$ (23)
For
$1\leq r\leq T^{1/4}$ (24)
we have $\log(1+\frac{r}{m})=\frac{r}{m}+O((r/m)^{2})$ and $e^{O({Tr^{2}/m^{2}})}=1+O({Tr^{2}/m^{2}})$. Applying this together with (23) into the right-hand side of (17) yields
@@
as $X\rightarrow\infty$. 222see e.g. https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
@@
For $\|x_{m}\|>Z^{-1}$ we use $\frac{1}{1+Z\|x_{m}\|}\leq\frac{1}{Z\|x_{m}\|}.$ Partition $(0,1/2]$ dyadically: $2^{-j-1}<\|x_{m}\|\leq 2^{-j}$ where $0\leq j\leq\log_{2}Z.$ For such $m$, $\frac{1}{1+Z\|x_{m}\|}\ll\frac{2^{j}}{Z}$. Since $x_{m}$ is monotone in $m$, the number of $m$ in each dyadic interval is $\ll T\cdot 2^{-j}.$ Hence the contribution of each $j$ is $\ll T\cdot 2^{-j}\cdot\frac{2^{j}}{Z}=\frac{T}{Z}.$ Summing over $0\leq j\leq\log_{2}Z$ yields